OpenStudy (anonymous):

What is the domain for the following function? y=2x/x-3 A. {x≠0} B. {x≠-3} C. {x≠3} D. {all real numbers}

jimthompson5910 (jim_thompson5910):

You can't divide by zero. So x-3 can't equal zero. So this means that x can't equal zero. So the answer is choice C

jimthompson5910 (jim_thompson5910):

oops I meant x can't equal 3 (since x-3=3-3=0)

OpenStudy (anonymous):

thank you!